A sonometer wire of l vibrates at a frequency of 350hz when the length of the wire is increased by 15cm the wire vibrates at 280hz under constant tension determine l 


1
Expert's answer
2021-07-25T09:16:37-0400
"f_1=\frac{v}{2L}\\f_2=\frac{v}{2(L+0.15)}\\\frac{350}{280}=\frac{L+0.15}{(L)}\\L=0.6\ m"